Grade 12Integral Calculus

integrate the following integral
∫ (e^x + 1)^2 dx

taniska
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int (e^x+1)^2dx
I = \int (e^{2x}+1+2e^x)dx
I = \int e^{2x}dx+\int dx+2\int e^xdx
I = \frac{e^{2x}}{2}+\int dx+2\int e^xdx
I = \frac{e^{2x}}{2}+x+2\int e^xdx
I = \frac{e^{2x}}{2}+x+2e^x+c